OpenStudy (wilder.monday):

Expand the following using the binomial theorem and Pascal’s triangle. 3.(2x + 3)^5

OpenStudy (wilder.monday):

@ganeshie8

OpenStudy (wilder.monday):

I've done up to a certain point but I don't know how to go on.

ganeshie8 (ganeshie8):

pick the 5th row for coefficients http://ptri1.tripod.com/ptreal1r.gif

OpenStudy (wilder.monday):

I know but I thought the 2 in 2x was the coefficient which is why I'm confused.

ganeshie8 (ganeshie8):

\[\begin{align} (2x+3)^5 &= \square (2x)^5(3)^0 + \square (2x)^4(3)^1 +\square (2x)^3(3)^2 +\square (2x)^2(3)^3 \\~\\&+\square (2x)^1(3)^4 +\square (2x)^0(3)^5 \end{align}\]

ganeshie8 (ganeshie8):

fill in the little squares with the coefficients from 5th row of pascal triangle

OpenStudy (wilder.monday):

okay and then just multiply the numbers by 2x?

ganeshie8 (ganeshie8):

kindof : \[\begin{align} (2x+3)^5 &= \bbox[3px, border:2px solid black]{1} (2x)^5(3)^0 + \bbox[3px, border:2px solid black]{5} (2x)^4(3)^1 + \bbox[3px, border:2px solid black]{10} (2x)^3(3)^2 + \bbox[3px, border:2px solid black]{10} (2x)^2(3)^3 \\~\\&+ \bbox[3px, border:2px solid black]{5} (2x)^1(3)^4 + \bbox[3px, border:2px solid black]{1} (2x)^0(3)^5 \end{align}\]

ganeshie8 (ganeshie8):

simplify each term

ganeshie8 (ganeshie8):

whats the valie of (2x)^5 ?

OpenStudy (wilder.monday):

uh...32x ?

ganeshie8 (ganeshie8):

the exponent applies to both the factors (2x)^5 = 2^5*x^5 = 32x^5

OpenStudy (wilder.monday):

so just keep doing that with all of them...

ganeshie8 (ganeshie8):

Yep

OpenStudy (wilder.monday):

so how would I do it for 10x^4 because that's 10,000x^4 ?

ganeshie8 (ganeshie8):

are you refering to below term ?\[\large 5(2x)^4(3)^1\]

OpenStudy (wilder.monday):

yes :)

ganeshie8 (ganeshie8):

first work the parenthesis : \[\large 5(2^4x^4)(3^1)\]

ganeshie8 (ganeshie8):

simplifying gives \[\large 5(16x^4)(3)\] furthher simplifying you get \[\large 240x^4\]

ganeshie8 (ganeshie8):

go ahead, try simplifying remaining terms. im sure you will be able to work them :)

OpenStudy (wilder.monday):

ohhhhh I get it now thanks

ganeshie8 (ganeshie8):

nice, so whats the final answer ?

OpenStudy (wilder.monday):

\[32x ^{5} + 240x ^{4} + 720x ^{3} + 1080x ^{2} + 810x + 243\]

ganeshie8 (ganeshie8):

Perfect! you may double check your answer wid wolfram http://www.wolframalpha.com/input/?i=expand+%282x+%2B+3%29%5E5

OpenStudy (wilder.monday):

thank you so much!! can you help with one more? or just the first part because I think its the same process but I want to make sure I get it right?

ganeshie8 (ganeshie8):

wil try, please open a new q :)